Which is the solution to the system?

x+5y = -15
y = 1/5x + 1

Below is how I solved it:
x+5(1/5x+1)=-15
x+(x+5)=-15
x^2+5=-15
x^2+5-5=-15-5
x^2=-20

Is this correct ?



Answer by Alan3354(69443)

x+5y = -15
y = 1/5x + 1

Below is how I solved it:
x+5(1/5x+1)=-15
x+(x+5)=-15
x^2+5=-15
x^2+5-5=-15-5
x^2=-20

Is this correct ?
-----------------
x+5(1/5x+1)=-15 Sub for y is good.
x+(x+5)=-15 This is correct.
x^2+5=-15 *** Don't multiply, add.
=============
2x+5 = -15
2x = -20
x = -10
=============
x^2+5-5=-15-5
x^2=-20